[figure: 9.7 cm]

COQBLIN-SCHRIEFFER IMPURITY WITH LOCAL EXCHANGE COUPLING: A PERTURBATIVE APPROACH

Piotr Stefański Institute of Molecular Physics of the Polish Academy of Sciences Smoluchowskiego 17, 60-179 Poznań, POLAND

The Coqblin-Schrieffer (C-S) one-impurity model with local exchange (LE) interaction is studied within the diagrammatic approach. The RKKY-type LE interaction between f electrons of the impurity and the additional non-hybridizing channel of conduction electron l=0 partial waves is treated in the molecular field approximation. The perturbation expansion resummation for the C-S hybridization interaction vertex part is carried out in the ladder approximation yielding the formula for the Kondo temperature in the presence of the LE interaction. It is found that the Kondo temperature T_{K} decreases whith increasing LE, and the temperature divergence of the susceptibility at T_{K} is shifted towards T=0. For some critical strength of the LE interaction the susceptibility approaches Curie-type dependence as for uncompensated impurity magnetic moment. A relation to the ’Kondo disorder’ model, which leads to the non-Fermi-liquid behavior, is discussed.

[figure: 13.4 cm]

Subject category :

1. Correlated Electrons and High Temperature Superconductors

Presentation mode : poster

Corresponding author : Piotr Stefański

Address for correspondence : Institute of Molecular Physics of the Polish Academy of Sciences Smoluchowskiego 17, 60-179 Poznań, POLAND

Email address : piotrs@ifmpan.poznan.pl